Successive differentiation ofTan inverse of (x/a) jdjfjfkdudu

Grade:12th pass

1 Answers

Vikas TU
6 years ago
 We realize that d/dx(tan-1x)=1/1+x2 
Additionally tan-1(x/a)=1/1+(x/a)2 
=1/1+x2/a2 
=a2/a2+x2 
Hence this is the differentiation.